For two independent events A and B, if P(A) = 2.0 and P(B) = 4.0, then 

(A∩ B) =0 .6

Check either true or false with shirt proof or counter example


1
Expert's answer
2021-04-14T14:47:47-0400

Events A and B are independent if


"P(A\cap B)=P(A)\times P(B)"

If A and B are not independent then they are dependent.


In this case:

P(A)= 2.0

P(B)=4.0

and P("A\cap B" )= 0.6

Now, applying condition

P(A)"\times P(B)" = 2.0"\times" 4.0= 8.0 which is not equal to "P(A\cap B)"

So, Events A and B are dependent events.